# Machine Learning and the Physical Sciences

### v1

@@ 1 Introduction
NOvA [2] is a long-baseline neutrino oscillation experiment in which neutrinos from Fermilab’s NuMI beam interact in both the near detector and larger far detector spaced 810km apart. The detectors consist of long PVC extrusions, referred to as "cells", filled with liquid scintillator. These cells are arranged in planes such that the cells’ orientations alternate horizontally and vertically between neighboring planes (Figure 1(a)). When neutrinos interact within the detector, the resulting particles, e.g. pions, photons, protons, electrons and muons, produce scintillation light as they travel through the detector and appear as tracks or showers referred to as "prongs" (Figure 1(b)). The task at hand is then to perform instance segmentation over these prongs as well as to classify each detection into its corresponding particle type.
@@
The design of the NOvA detectors only allows images to be produced over two planes: the XZ plane (top view) and YZ plane (side view), in which each view only consists of vertical or horizontal cells, respectively (Figure 1(a)). Therefore, the Y position of a hit in a vertical plane cannot be known. Existing methods in use at NOvA approach this by running two independent CNNs to embed each image [1, 8, 4] or by using each view as a channel to an image [9]. We can extend the framework of point transformers by using heterogeneous attention [7]. In this paper we leverage point set transformers, along with implementing a novel method to allow information to flow between both views of the dataset, allowing for a model that is able to draw information from both views. The model proposed is known as a heterogeneous point set transformer (HPST).
@@ 2.2 Heterogeneous attention
We use the inter-view distance to build a nearest neighbors graph, then, for each point we calculate the query $Q^{(i,j^{\prime}\to j)}_{k}$, i.e., the query on point $k$ from view $j^{\prime}$ to view $j$ on sample $i$, and then for each of its neighbors $k^{\prime}$ we calculate both $K^{(i,j^{\prime}\to j)}_{k^{\prime}}$ and $V^{(i,j^{\prime}\to j)}_{k^{\prime}}$, that is, the key and values on point $k^{\prime}$ from view $j^{\prime}$ to view $j$ on sample $i$. Using these, we can calculate $w^{(i,j^{\prime}\to j)}_{kk^{\prime}}=Q_{k}^{(i,j^{\prime}\to j)T}K^{(i,j^{\prime}\to j)}_{k^{\prime}}$. RPE is not used between different views. This weight is then normalized using a softmax operation over its neighbors and then used in a weighted sum to calculate the output of the attention module,
@@ 3.1 Experiments
Three models were trained and evaluated: a Mask-RCNN [6] based model, GAT [11], and our heterogeneous point set transformer. We performed a hyperparameter search: using the number of neighbor connections (4, 8), the number of stages of the neural network (2, 3, 4), the size of the embeddings inside the neural network (128, 256, 512), and the learning rate (between 1e-4 and 1e-1). The training and testing was done on a server using an Intel(R) Xeon(R) CPU E5-2640 v4 @ 2.40GHz, 503G of RAM, and 4xNVIDIA Titan V. The hyperparameter sweep was peformed over 8 epochs, using 1% of the dataset. The model with the best accuracy on the segmentation’s class labels was selected as the one with the best hyperparameters. The resulting models with the best hyperparameters were trained for 24 epochs.
@@ 4 Limitations and Conclusions
While the claims of memory efficiency will generally hold true, for different datasets this might not be the case. The representation of a sparse matrix is more efficient than a dense matrix until a certain point, where the storage of the coordinates becomes bigger than just storing a dense matrix. Point set operations can also greatly increase in complexity as the number of points grows, resulting in a much slower algorithm. However, these are not the regimes found in the the NOvA experiment. HPSTs strike a balance between memory usage, time, and performance that is fitting for its application.

### v2

@@ 1 Introduction
[…]
@@
The design of the NOvA detectors only allows images to be produced over two planes: the XZ plane (top view) and YZ plane (side view), in which each view only consists of vertical or horizontal cells, respectively (Figure 1(a)). Therefore, the Y position of a hit in a vertical plane fundamentally cannot be known. The existing method for the instance segmentation task approaches this by applying the fuzzy k-means clustering algorithm on each view independently followed by matching instances across the two views [4]. Similarly for the particle classification task, existing methods use two independent CNNs to embed each image [2, 10, 6] or use each view as a channel to an image [11]. We can extend the framework of point transformers by using heterogeneous attention [9]. In this paper we leverage point set transformers, along with implementing a novel method to allow information to flow between both views of the dataset, allowing for a model that is able to draw information from both views. The model proposed is known as a heterogeneous point set transformer (HPST).
@@ 2.2 Heterogeneous attention
In order to calculate attention, we need to encode each point into a set of keys, queries and values, each of these is done by taking the input point $x^{(i,j)}_{k}$ and multiplying it by a linear layer as such: $Q^{(i,j^{\prime}\to j)}_{k}=W^{(j^{\prime}\to j)}_{Q}x^{(i,j)}_{k}$, $K^{(i,j^{\prime}\to j)}_{k}=W^{(j^{\prime}\to j)}_{K}x^{(i,j)}_{k}$, $V^{(i,j^{\prime}\to j)}_{k}=W^{(j^{\prime}\to j)}_{V}x^{(i,j)}_{k}$.
We use the inter-view distance to build a nearest neighbors graph, then, for each point we calculate the query $Q^{(i,j^{\prime}\to j)}_{k}$, i.e., the query on point $k$ from view $j^{\prime}$ to view $j$ on sample $i$, and then for each of its neighbors $k^{\prime}$ we calculate both $K^{(i,j^{\prime}\to j)}_{k^{\prime}}$ and $V^{(i,j^{\prime}\to j)}_{k^{\prime}}$, that is, the key and values on point $k^{\prime}$ from view $j^{\prime}$ to view $j$ on sample $i$. Using these, we can calculate $w^{(i,j^{\prime}\to j)}_{kk^{\prime}}=Q_{k}^{(i,j^{\prime}\to j)T}K^{(i,j^{\prime}\to j)}_{k^{\prime}}$. RPE (relative positional encoding, $\mathrm{RPE}(x,y)=W(x-y)$, where $W$ is a learnable linear layer as used in [14]) is not used between different views. This weight is then normalized using a softmax operation over its neighbors and then used in a weighted sum to calculate the output of the attention module,
@@ 3.1 Experiments
Three models were trained and evaluated: a Mask-RCNN [8] based model, GAT [13], and our heterogeneous point set transformer. We performed a hyperparameter search: using the number of neighbor connections (4, 8), the number of stages of the neural network (2, 3, 4), the size of the embeddings inside the neural network (128, 256, 512), and the learning rate (between 1e-4 and 1e-1). The training and testing was done on a server using an Intel(R) Xeon(R) CPU E5-2640 v4 @ 2.40GHz, 503G of RAM, and 4xNVIDIA Titan V. The hyperparameter sweep was peformed over 8 epochs, using 1% of the dataset. The model with the best accuracy on the segmentation’s class labels was selected as the one with the best hyperparameters. The resulting models with the best hyperparameters were trained for 24 epochs. The code for these experiments can be found at https://github.com/erobl/hpst.
@@ 4 Limitations and Conclusions
While the claims of memory efficiency will generally hold true, for different datasets this might not be the case. The representation of a sparse matrix is more efficient than a dense matrix until a certain point, where the storage of the coordinates becomes bigger than just storing a dense matrix. Point set operations can also greatly increase in complexity as the number of points grows, resulting in a much slower algorithm. However, these are not the regimes found in the the NOvA experiment. HPSTs strike a balance between memory usage, time, and performance that is fitting for its application, they are able to train a stronger model with the same amount of parameters due to sharing information between the views, which in turn boosts the performance of the model.